# Making FETCH! Happen: Finding Emergent Dog Whistles Through Common Habitats

### v1

@@ Abstract
Dog whistles are coded expressions with dual meanings: one intended for the general public (outgroup) and another that conveys a specific message to an intended audience (ingroup). Often, these expressions are used to convey controversial political opinions while maintaining plausible deniability and slip by content moderation filters. Identification of dog whistles relies on curated lexicons, which have trouble keeping up to date. We introduce FETCH!, a task for finding novel dog whistles in massive social media corpora. We find that state-of-the-art systems fail to achieve meaningful results across three distinct social media case studies. We present EarShot, a novel system that combines the strengths of vector databases and Large Language Models (LLMs) to efficiently and effectively identify new dog whistles. [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/FETCH-Dog-Whistle
@@ 1 Introduction
When a person writes dual citizens it may seem like an innocuous phrase, but many understand this as an antisemitic reference. The use of the phrase builds on centuries old accusations that question the loyalty of Jewish people towards state governments, which in recent decades has become a claim that all American Jews are double citizens of Israel and the United States.
This phrase is a dog whistle, which is defined as an “expression that send one message to an outgroup while at the same time sending a second (often taboo, controversial, or inflammatory) message to an ingroup Henderson and McCready (2018).” Dog whistles are a common mainstay in today’s turbulent political climate. For example, they prominently featured in the 2016 United States Presidential Election both on social media and in person Tilley et al. (2020); Drakulich et al. (2020).
@@
As dog whistles are a common way to transmit hate speech to a wide audience both digitally and offline, researchers have started to develop methods to detect them. However, dog whistles are notoriously difficult to find as they often slip by content moderation, toxicity, and hate speech filters at a high rate because of their benign second meaning Mendelsohn et al. (2023). Dual citizenship isn’t an issue, but “dual citizen” in the above context has a specific meaning.
@@
We consider the task of dog whistle discovery, a novel task where systems identify emerging or previously unknown dog whistles. To support this effort, we create FETCH! or Finding Emergent Dog Whistles Through Common Habitats, a benchmark designed to evaluate models’ ability to detect novel dog whistles across diverse settings. Using the FETCH! benchmark, we evaluate three state of the art methods across three different social media case studies. Our evaluation reveals that these methods perform poorly, achieving an F-score of less than 5% across all case studies.
We introduce our own method: EarShot. This method finds new dog whistles by leveraging the power of sentence embeddings and Large Language Models (LLMs). EarShot demonstrates significant improvements, with F-score increases ranging from $2$ to $20$ points across the social media case studies. We make the following key contributions:
  • Development of FETCH! benchmark for dog whistle discovery;
  • Rigorous evaluation of multiple state-of-the-art methods on FETCH!;
  • Creation of EarShot, a novel architecture achieving SoTA performance on FETCH!.
@@ 2.4 Dog Whistle Detection
Relatively few studies have addressed the detection of dog whistles. Among these, three notable works focus specifically on Swedish dog whistles, employing both embedding-based methods and analyzing semantic shifts Hertzberg et al. (2022); Boholm and Sayeed (2023); Boholm et al. (2024). In English, only two studies have focused on detecting dog whistles. In Mendelsohn et al. (2023), they curated set of English dog whistles to assess both large language models’ (LLMs) ability to surface these terms and the impact on hate speech detectors. However, they only do a small analysis with a handful of examples to see if LLMs can detect novel dog whistles. In Kruk et al. (2024), they introduced the largest dataset of posts containing dog whistles, filtered from sources such as Reddit and the Congressional Record by LLMs. In addition to the creation of the dataset, they conducted a preliminary test of LLMs’ capabilities in detecting novel dog whistles.
@@ 3.1.1 Synthetic (Reddit)
Our first scenario is an idealized setting where every post contains an associated dog whistle. We use the Silent Signals dataset Kruk et al. (2024), which includes 16,000 posts scraped from Reddit. Each post in the dataset contains a dog whistle identified by GPT-4 OpenAI et al. (2024).
@@ 3.1.2 Balanced (Gab)
The balanced scenario represents a context with a higher-than-average prevalence of dog whistles with large amount of posts. For this scenario, we chose to use data from Gab, an alt-right alternative to Twitter where dog whistles and other forms of hate speech are more frequent. The data was collected over a short period using a web scraper that sequentially downloaded post IDs. This process yielded approximately 300,000 posts.
@@ 3.2 Seed Dog Whistles
We use the dog whistle dataset released by Mendelsohn et al. (2023) as our ground truth. The dataset is a hand-curated taxonomy of around 340 root dog whistles from community wikis and academic sources. Each root dog whistle comes with the usage of the word, the group it is targeting, the group that often uses it, definitions, and other surface forms of the root dog whistle. We use these dog whistles for the Realistic and Balanced scenarios as the Synthetic scenario has its own set of dog whistles. Table 1 lists a summary of the number of dog whistles found in each scenario. The full list appears in Appendix A.
@@
For each corpora, we divide the dog whistles into train/test splits (20% / 80%), creating a set of seed dog whistles and dog whistles to find. We split the dog whistles using stratified sampling based on how the length of phrase in $n$-grams. We do a stratified split so as not to unfairly punish methods that can identify larger $n$-grams. After splitting, we add the surface forms as well to each split in order to cover all variants of the dog whistles.
@@ 3.3 Metrics for Evaluation
We use precision, recall and F-score to measure the performance of different methods for our task. Precision is compute as normal, but we modify Recall to normalize by only the positive terms that exist in the corpus (Data Potential Recall, DPR). Due to the multiple surface forms of dog whistles provided in the dataset, all metrics count a positive match for a single root dog whistle as returning an exact match of one of the surface forms or the root itself.
We report $F_{0.5}$ scores using DPR and Precision instead of $F_{1}$ to preference methods with a high precision to reduce the load on a human reviewer.
@@ 4.1.1 Word2Vec/Phrase2Vec
To create the training data for the Word2Vec/Phrase2Vec Mikolov et al. (2013), we preprocessed the text to remove all stop words and lemmatize the words using littlebird DeLucia (2020). After this preprocessing step, we replaced URLs, hashtags, retweets, and mentions with special symbols like @USER and HTTPURL. We also replaced emojis with their official English aliases using the emoji package Kim and Wurster (2024). Finally, we tokenized the result with the NLTK Tweet Tokenizer Bird and Loper (2004).
To construct the Phrase2Vec models, we utilized the Gensim Phraser module Rehurek and Sojka (2011) to merge common $n$-gram phrases in the original dataset, generating $2$-gram and $3$-gram versions of the models. We only train up to $3$-gram models as a majority of dog whistles in our dataset are $3$-grams and smaller. We then trained models using Gensim with a maximum vocabulary size of 500,000 and a context window of 5 with a vector dimension of 100 for 10 epochs
@@ 4.1.2 MLM
Euphemism detection relies heavily on context, but static embeddings from Word2Vec/Phrase2Vec fail to capture this. To address this, Zhu et al. (2021) introduces a model leveraging contextual embeddings from Masked Language Models like BERT. The MLM system first samples 2000 sentences from the input dataset containing the seed dog whistles. We sampled 2000 sentences as it is default parameter provided. For each sampled sentence, it masks out the seed dog whistle in that sentence and uses a Masked Language Model to fill in the mask by creating a distribution over possible replacements and then returning the top-$k$ most likely. We test this method using Bernice DeLucia et al. (2022) and BERTweet Nguyen et al. (2020), two MLMs that have already been adapted to social media data. We test the method at thresholds of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
@@ 4.1.3 EPD
The Euphemistic Phrase Detector (EPD) Zhu and Bhat (2021) builds upon the MLM model, addressing its limitation of restricting predictions to a single token from the tokenizer. To solve this issue, the EPD model generates possible phrases using Autophrase Shang et al. (2017), a statistical phrase extractor to choose important phrases from a corpus. Next, it trains a unigram Word2Vec model Mikolov et al. (2013) on the same corpus. It uses this model to filter down the phrases by selecting the top 1000 most similar phrases from AutoPhrase to the seed dog whistles. Finally, similar to MLM method, it samples 2000 sentences from the input dataset that contain the seed dog whistles. For each sampled sentence, it replaces the seed dog whistle with each phrase kept from the previous stage and uses a Masked Language Model to calculate the probability of the sentence with the replaced phrase. They then sort the sentences via their probability and select the top-$k$ sentences and return the corresponding inserted phrase as predictions. We test this method using SpanBERT Joshi et al. (2020) as our Masked Language Model. We test the method at increasing thresholds of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
@@ 4.1.4 EarShot
We propose a new method called EarShot. EarShot consists of three stages.
@@
In the second stage, we obtain the vectors for the posts containing the seed dog whistles and find the closest vector to each post vector that are not the post themselves. By identifying closet neighbors, we aim to capture posts that are likely related in meaning, sentiment, or intent, even if they do not share exact word matches.
After this step, there are two different pipelines we try: DIRECT and PREDICT, labeled (1) and (2) in Figure 2 respectfully.
For the DIRECT pipeline, we pass all the posts to a LLM to directly extract the dog whistles from the posts provided from the vector lookup. Similarly to Kruk et al. (2024), the model is prompted to list all dog whistles in a structured JSON format. The prompt is shown in Table 7 in Appendix B. We evaluate three LLMs: meta-llama/Meta-Llama-3.1-8B-Instruct (LLaMa 8B) Dubey et al. (2024), mistralai/Mistral-7B-Instruct-v0.3 (Mistral 7B) Jiang et al. (2023), and meta-llama/Llama-2-13b-chat-hf (LLaMa 13B) Touvron et al. (2023). We choose to test LLaMa and Mistral as they some of the best performing open source LLMs.
@@
For the LLM based filtering, we prompt the model to reply yes or no if there is a dog whistle in the post. We structure our prompt similarly to Silent Signals Kruk et al. (2024), shown in Table 7 in Appendix B. We keep all the posts that the model explicitly says yes to using a substring search. We call this filtering method LLM-PREDICT. We test the same three models used in the DIRECT pipeline.
After the filtering step, we pass the posts to a keyword extractors and return the top-$k$ predictions of them. We use five different popular keyword extraction algorithms: KeyBERT Grootendorst (2020), RAKE Rose et al. (2010), YAKE, Campos et al. (2018), TextRank Mihalcea and Tarau (2004), and TF-IDF. If the keyword extractor requires a specific range of $n$-grams to return, we test ranges of $1$-gram, 11112222-grams, and 11113333-grams. We report the performance of the method at the same thresholds in MLM and EPD.
@@ 5.1 Word2Vec/Phrase2Vec
We summarize the best models for Word2Vec and Phrase2Vec in Table 2. Overall, unigram models do the best across all three datasets both in terms of Precision and $F_{0.5}$. However, unigram Word2Vec models fail when it comes to recall often getting $2$-$13$ points less in recall as they cannot capture more complex phrases.
@@
Finally, most models do the best on the first or second level of prediction. This conservative prediction strategy decreases the usefulness of the model as it only predicts a small amount of new dog whistles. Figures displaying the choice of threshold is provided in Appendix C. Overall, unigram Word2Vec models do the best in this task but are unable to capture more complex phrases due to the limitations in the model and their conservative prediction strategy.
@@ 5.2 MLM and EPD
Table 3 presents the top-performing MLM and EPD models for each scenario. Overall, MLM and EPD models perform extremely poorly across all metrics. They obtain less than $1$ point in F-score, very poor precision around $1$-$2$ points, and a modest DPR around $1$-$9$ points. For the MLM models, we attribute this lackluster performance due to limitation of the tokenizer of the models as MLM models are restricted to predicting single tokens learned from their training corpora. As for the EPD model, we believe that the filtering step using a unigram Word2Vec model leaves out complex phrases.
Looking deeper, most of the models only predict 50 words being extremely conservative with their choices which does lead to better precision, mimicking the same failure mode in Word2Vec and Phrase2Vec. Additional figures detailing the choice of threshold are provided in Appendix D.
As for model choice, EPD does the best in Synthetic and Realistic but does not work in Balanced. We hypothesize that SpanBERT is not as adapted to a domain as different as Balanced reducing its effectiveness. Overall, MLM and EPD models have lackluster performance and are beat by the simpler Word2Vec and Phrase2Vec models.
@@ 5.3.1 PREDICT
Table 4: Best Results for FETCH LLM and BERT PREDICT based on $F_{0.5}$. Prec is Precision. DPR is Data Potential Recall. Best scores for across each dataset are bolded
In Table 4, we choose the best LLM and BERT filtering model for each dataset based on the $F_{0.5}$ of the system. EarShot-Predict obtains high precision of around $9$-$20$ points, beating Word2Vec, Phrase2Vec, MLM, and EPD. However, its DPR suffers obtaining around $1$-$13$ points. While it beats MLM and EPD in terms of DPR, Word2Vec and Phrase2Vec beat PREDICT on Balanced and Synthetic and tie it on Realistic. This low recall stems from the nature of keyword extraction models, which aim to identify important words rather than specifically targeting dog whistles themselves.
To achieve this high precision, the system makes the conservative choice of predicting on average less than 1000 words and only chooses $1-$grams. While its prediction abilities are still restricted, this conservative prediction threshold is an improvement over the other systems as it maintains high precision at larger thresholds, making it more competitive.
Looking towards the choices of keyword extractors, KeyBERT does the best on smaller datasets like Balanced and Synthetic and YAKE and RAKE do better on larger datasets like Realistic. For the LLM model choices, LLaMa 13B does the best across Balanced and Realistic while LLaMa 8B does better on Synthetic. As for the BERT model choices, RoBERTa R4 does best across Balanced and Synthetic while HateXplain BERT does well on Realistic. Figures detailing choices of keyword extraction models, filtering models, and thresholds can be found in Appendix E.
Comparing LLMs vs BERT for filtering, BERT-based models outperform LLMs across all three datasets by $0.1$ $0.5$ points. Even with all the extra parameters, task specific models are able to match the performance of larger general purpose LLMs for this task.

### v2

@@ Abstract
Dog whistles are coded expressions with dual meanings: one intended for the general public (outgroup) and another that conveys a specific message to an intended audience (ingroup). Often, these expressions are used to convey controversial political opinions while maintaining plausible deniability and slip by content moderation filters. Identification of dog whistles relies on curated lexicons, which have trouble keeping up to date. We introduce FETCH!, a task for finding novel dog whistles in massive social media corpora. We find that state-of-the-art systems fail to achieve meaningful results across three distinct social media case studies. We present EarShot, a strong baseline system that combines the strengths of vector databases and Large Language Models (LLMs) to efficiently and effectively identify new dog whistles.  [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/FETCH-Dog-Whistle
@@ 1 Introduction
When a person writes dual citizens it may seem like an innocuous phrase, but many understand this as an antisemitic reference. The use of the phrase builds on centuries old questions of the loyalty of Jewish people towards governments, which recently has become a claim that all American Jews are double citizens of Israel and the United States.
This phrase is a dog whistle, which is defined as an “expression that sends one message to an outgroup while at the same time sending a second (often taboo, controversial, or inflammatory) message to an ingroup Henderson and McCready (2018).” Dog whistles are a common mainstay in today’s turbulent political climate. For example, they prominently featured in the 2016 United States Presidential Election both on social media and in person Tilley et al. (2020); Drakulich et al. (2020).
@@
As dog whistles are a common way to transmit hate speech to a wide audience both digitally and offline, researchers have started to develop methods to detect them. However, dog whistles are notoriously difficult to find as they often slip by content moderation, toxicity, and hate speech filters at a high rate because of their benign second meaning Mendelsohn et al. (2023). Dual citizenship isn’t an issue, but “dual citizen” in the above context has a specific meaning. More examples with explanations of the dog whistles’ meaning and their difficulty are provided in Appendix B.
@@
We consider the task of dog whistle discovery, a novel task where systems identify emerging or previously unknown dog whistles. To support this effort, we create FETCH! or Finding Emergent Dog Whistles Through Common Habitats, a benchmark designed to evaluate models’ ability to detect emerging or novel dog whistles across diverse settings. Using the FETCH! benchmark, we evaluate three state of the art methods across three different social media case studies. Our evaluation reveals that these methods perform poorly, achieving an F-score of less than 5% across all case studies.
We introduce our own method as a strong baseline: EarShot. This method finds new dog whistles by leveraging the power of sentence embeddings and Large Language Models (LLMs). EarShot demonstrates significant improvements, with F-score increases ranging from $2$ to $20$ points across the social media case studies. We make the following key contributions: Development of FETCH! benchmark for dog whistle discovery; Rigorous evaluation of multiple state-of-the-art methods on FETCH!; Creation of EarShot, a strong baseline architecture achieving SoTA performance on FETCH!;
@@ 2.4 Dog Whistle Detection
Relatively few studies have addressed the detection of dog whistles. Among these, three notable works focus specifically on Swedish dog whistles, employing both embedding-based methods Boholm et al. (2024); Hertzberg et al. (2022) and analyzing semantic shifts Boholm and Sayeed (2023) In English, only two studies have focused on detecting dog whistles. In the first paper, they curated a set of English dog whistles to assess both large language models’ (LLMs) ability to surface these terms and the impact on hate speech detectors. However, they only do a small analysis with a handful of positive examples to see if LLMs can detect novel dog whistles Mendelsohn et al. (2023). In the second paper, they introduced the largest dataset of posts containing dog whistles, filtered from sources such as Reddit and the Congressional Record by LLMs. In addition to the creation of the dataset, they conducted a small preliminary test of LLMs’ capabilities in detecting novel dog whistles Kruk et al. (2024).
@@ 3.1.1 Synthetic (Reddit)
Our first scenario is an idealized setting where every post contains an associated dog whistle. We use the Silent Signals dataset Kruk et al. (2024), which includes 16,000 posts scraped from Reddit from 45 controversial subreddits. Each post in the dataset contains a dog whistle identified by GPT-4 OpenAI et al. (2024).
@@ 3.1.2 Balanced (Gab)
The balanced scenario represents a context with a higher-than-average prevalence of dog whistles with large amount of posts. For this scenario, we chose to use data from Gab, an alt-right alternative to Twitter where dog whistles and other forms of hate speech are more frequent. The data was collected over a short period using a web scraper that sequentially downloaded post IDs yielding approximately 300,000 posts.
@@ 3.2 Seed Dog Whistles
We use the dog whistle dataset released by Mendelsohn et al. (2023) as our ground truth. The dataset is a hand-curated taxonomy of around 340 root dog whistles from community wikis and academic sources. Each root dog whistle comes other surface forms of the root dog whistle. We use these dog whistles for the Realistic and Balanced scenarios as the Synthetic scenario has its own set of dog whistles. Table 1 lists a summary of the number of dog whistles found in each scenario. The full list appears in Appendix A.
@@
For each corpora, we divide the dog whistles that are able to be found into train/test splits (20% / 80%), creating a set of seed dog whistles to use to find the rest of the dog whistles. We split the dog whistles using stratified sampling based on the length of phrase in $n$-grams. We do a stratified split so as not to unfairly punish methods that can identify larger $n$-grams. After splitting, we add the surface forms as well to each split in order to cover all variants of the dog whistles. We split on $n$-grams over different metadata provided in the dog whistle list like the group being targeted or the user of the dog whistle as the Synthetic scenario does not provide this information and it is more realistic scenario to have a set of randomly chosen dog whistles rather a well distributed set of dog whistles according to the metadata.
@@ 3.3 Metrics for Evaluation
We use precision, recall and F-score to measure the performance of different methods for our task. We opt to use precision, recall, and F-score without a threshold as it is a more general metric as some methods do not rank their predictions. In addition, we do not create metrics that normalize for the maximum length of the words output as in testing, models with smaller length outputs have inflated scores.
Due to the multiple surface forms of dog whistles provided in the dataset, all metrics count a positive match for a single root dog whistle as returning an exact match of one of the surface forms or the root itself.
Precision is computed as normal, but we modify recall to normalize by only the positive terms that exist in the corpus (Data Potential Recall, DPR). We report $F_{0.5}$ scores using DPR and Precision instead of $F_{1}$ to preference methods with a high precision to reduce the load on a human reviewer.
@@ 4.1.1 Word2Vec/Phrase2Vec
To create the training data for the Word2Vec/Phrase2Vec Mikolov et al. (2013), we preprocessed the corpus by removing all stop words and lemmatizing the rest using littlebird DeLucia (2020). After this preprocessing step, we replaced URLs, hashtags, retweets, and mentions with special symbols like @USER and HTTPURL. We also replaced emojis with their official English aliases using the emoji package Kim and Wurster (2024). Finally, we tokenized the result with the NLTK Tweet Tokenizer Bird and Loper (2004).
To construct the Phrase2Vec models, we utilized the Gensim Phraser module Rehurek and Sojka (2011) to merge common $n$-gram phrases in the original corpus, generating $2$-gram and $3$-gram versions of the models. We only train up to $3$-gram models as a majority of dog whistles in our dataset are $3$-grams and smaller. We then trained models using Gensim with a maximum vocabulary size of 500,000 and a context window of 5 with a vector dimension of 100 for 10 epochs.
@@ 4.1.2 MLM
Euphemism detection relies heavily on context, but static embeddings from Word2Vec/Phrase2Vec fail to capture this. To address this, researchers leveraged contextual embeddings from Masked Language Models like BERT Zhu et al. (2021). The MLM system first samples 2000 sentences from the input dataset containing the seed dog whistles. We sampled 2000 sentences as it is the default parameter provided. For each sampled sentence, it masks out the seed dog whistle in that sentence and uses the Masked Language Model to fill in the mask by creating a distribution over possible replacements and then returning the top-$k$ most likely. We test this method using two MLM encoder models: Bernice DeLucia et al. (2022) and BERTweet Nguyen et al. (2020). Both these models have been adapted to social media data, providing a reasonable benchmark as pretraining custom models from scratch computationally expensive. We test the method at thresholds for $k$ of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
@@ 4.1.3 EPD
The Euphemistic Phrase Detector (EPD) Zhu and Bhat (2021) builds upon the MLM model by predicting phrases instead of single tokens. The EPD model uses Autophrase Shang et al. (2017), a statistical phrase extractor to extract important phrases from the corpus. Next, it trains a unigram Word2Vec model Mikolov et al. (2013) on the same corpus to filter down the phrases by selecting the top 1000 most similar phrases from AutoPhrase to the seed dog whistles. Finally, it samples 2000 sentences from the input dataset that contain the seed dog whistles. For each sampled sentence, it replaces the seed dog whistle with each predicted phrase from the previous stage and uses the Masked Language Model to calculate the probability of the sentence with the replaced phrase. They then sort the sentences via their probability and select the top-$k$ sentences and return the corresponding inserted phrase as predictions. We test this method using SpanBERT Joshi et al. (2020) as our Masked Language Model as it is the default model used. We test the method at increasing thresholds for $k$ of $\{50,100,200,400,800,1600,3200,6400,12800,\\ 25600\}$.
@@ 4.1.4 EarShot
We propose EarShot, a strong baseline architecture for the task, consisting of three stages.
@@
In the second stage, we obtain the vectors for the posts containing the seed dog whistles and find the closest vector to each post vector that are not the post themselves. By identifying closet neighbors, we aim to capture posts that are likely related in meaning, sentiment, or intent, even if they do not share exact word matches. After this step, we try two different methods: DIRECT and PREDICT, labeled (1) and (2) in Figure 2 respectfully.
For the DIRECT pipeline, we pass all the posts to a LLM to extract the dog whistles from the posts provided from the vector lookup. The model is prompted to list all dog whistles in a structured JSON format with a prompt from Kruk et al. (2024). The prompt is shown in Table 7 in Appendix C. We evaluate three LLMs: meta-llama/Meta-Llama-3.1-8B-Instruct (LLaMa 8B) Dubey et al. (2024), mistralai/Mistral-7B-Instruct-v0.3 (Mistral 7B) Jiang et al. (2023), and meta-llama/Llama-2-13b-chat-hf (LLaMa 13B) Touvron et al. (2023). We choose to test LLaMa and Mistral as they some of the best performing open source LLMs. We choose to use smaller open source models over larger closed source models as they are easier to reproduce and provide a strong baseline.
@@
For the LLM based filtering, we prompt the model to reply yes or no if there is a dog whistle in the post. We structure our prompt similarly to Silent Signals Kruk et al. (2024), shown in Table 7 in Appendix C. We keep all the posts that the model says yes to using a substring search. We call this filtering method LLM-PREDICT. We test the same three models used in the DIRECT pipeline.
After the filtering step, we pass the posts to a keyword extractors and return the top-$k$ predictions of them. We use five different popular keyword extraction algorithms: KeyBERT Grootendorst (2020), RAKE Rose et al. (2010), YAKE, Campos et al. (2018), TextRank Mihalcea and Tarau (2004), and TF-IDF Spärck Jones (1972). If the keyword extractor requires a specific range of $n$-grams to return, we test ranges of $1$-gram, 11112222-grams, and 11113333-grams. We report the performance of the method at the same thresholds in MLM and EPD.
@@ 5.1 Word2Vec/Phrase2Vec
We summarize the best models for Word2Vec and Phrase2Vec in Table 2. Unigram models do the best across all three datasets both in terms of Precision and $F_{0.5}$ but it comes at the cost of lower recall as they cannot capture more complex phrases.
@@
Most models perform best at the first or second prediction level. This conservative strategy limits the model’s usefulness, as it predicts only a small number of new dog whistles. Threshold choice figures are in Appendix D. Overall, unigram Word2Vec models excel in this task but struggle to capture more complex phrases due to model limitations and their conservative prediction approach.
@@ 5.2 MLM and EPD
Table 3 presents the top-performing MLM and EPD models for each scenario. Overall, MLM and EPD models perform extremely poorly across all metrics. They obtain less than $1$ point in F-score, very poor precision around $1$-$2$ points, and a modest DPR around $1$-$9$ points. For the MLM models, we attribute this lackluster performance to limitation of the tokenizer of the models as MLM models are restricted to predicting single tokens learned from their training corpora. As for the EPD model, we believe that the filtering step using a unigram Word2Vec model leaves out complex phrases.
Most of the models only predict 50 words mimicking failures in Word2Vec and Phrase2Vec. Additional figures detailing the choice of threshold are provided in Appendix E. Overall, MLM and EPD models lag behind the simpler embedding models.
@@ 5.3.1 PREDICT
Table 4: Best Results for FETCH LLM and BERT PREDICT based on $F_{0.5}$. Prec is Precision. DPR is Data Potential Recall. Best scores for each dataset are bolded
In Table 4, we choose the best LLM and BERT filtering model for each dataset based on the $F_{0.5}$ of the system. EarShot-PREDICT obtains higher precision than other systems with around $9$-$20$ points. However, its DPR suffers obtaining around $1$-$13$ points. While it beats MLM and EPD in terms of DPR, Word2Vec and Phrase2Vec beat PREDICT on Balanced and Synthetic and tie it on Realistic stemming from keyword extraction models, which aim to identify important words rather than specifically targeting dog whistles themselves.
To achieve this high precision, the system makes the conservative choice of predicting on average less than 1000 words and only chooses $1-$grams. While its prediction abilities are still restricted, this prediction threshold improves over the other systems as it maintains high precision.
Looking towards the choices of keyword extractors, KeyBERT does the best on smaller datasets like Balanced and Synthetic and YAKE and RAKE do better on larger datasets like Realistic. For the LLM model choices, LLaMa 13B does the best across Balanced and Realistic while LLaMa 8B does better on Synthetic. As for the BERT model choices, RoBERTa R4 does best across Balanced and Synthetic while HateXplain BERT does well on Realistic. Figures detailing choices of keyword extraction models, filtering models, and thresholds can be found in Appendix F.
BERT-based models outperform LLMs by 0.1–0.5 points across all three datasets. Despite their extra parameters, larger LLMs do not surpass task-specific models in this task.